Examine whether the equation, x^3-11x+9=0 has a real root in the interval, [-1,2]

1
Expert's answer
2022-02-01T12:23:35-0500

Solution:

"x^3-11x+9=0"

"f(x)=x^3-11x+9"

Interval: [-1,2]

"f(-1)=(-1)^3-11(-1)+9
\\=-1+11+9
\\=19
\\f(2)=(2)^3-11(2)+9
\\=8-22+9
\\=-5"

Now, "f(-1).f(2)=19(-5)=-95<0"

"\because f(-1).f(2)<0"

So, yes, a real root exists in the interval [-1,2]